OpenStudy (anonymous):

Solve the following system of equations by substitution: 4x − 6y = 28 8x − 4y = 24 x = 1, y = 4 x = −1, y = 4 x = 1, y = −4 x = −1, y = −4

OpenStudy (hybrik):

Multiply first equation by 2, 8x-12y=56 8x-12y=56 8x − 4y = 24 Subtract X, -8y=32, y = -4 plug in y into any of 2 equations

OpenStudy (hybrik):

4x-6(-4)=28, 4x+24=28, 4x=4, x = 1, {1,-4) = C

OpenStudy (anonymous):

Now I get it! Thank you so much!!!

OpenStudy (hybrik):

Your welcome :)